Week of 170501

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Tier-1 downtimes

Experiments may experience problems if two or more of their Tier-1 sites are inaccessible at the same time. Therefore Tier-1 sites should do their best to avoid scheduling a downtime classified as "outage" in a time slot overlapping with an "outage" downtime already declared by another Tier-1 site supporting the same VO(s). The following procedure is recommended:

  1. A Tier-1 should check the downtimes calendar to see if another Tier-1 has already an "outage" downtime in the desired time slot.
  2. If there is a conflict, another time slot should be chosen.
  3. In case stronger constraints cannot allow to choose another time slot, the Tier-1 will point out the existence of the conflict to the SCOD mailing list and at the next WLCG operations call, to discuss it with the representatives of the experiments involved and the other Tier-1.

As an additional precaution, the SCOD will check the downtimes calendar for Tier-1 "outage" downtime conflicts at least once during his/her shift, for the current and the following two weeks; in case a conflict is found, it will be discussed at the next operations call, or offline if at least one relevant experiment or site contact is absent.

Links to Tier-1 downtimes

ALICE ATLAS CMS LHCB
  BNL FNAL  

Monday: May Day holiday

  • The meeting will be held on Tuesday instead.

Tuesday

Attendance:

  • local: Kate (WLCG, DB), Julia (WLCG), Gavin (computing), Vincent (security), Andrea (MW)
  • remote: Onno (NL-T1), FaHui (ASGC), Luca L (CNAF), Kyle (OSG), Di Qing (TRIUMF), Dave M (FNAL), Christoph W (CMS), Xin Zhao (BNL), Nurcan (ATLAS), John (RAL), Pepe (PIC)

Experiments round table:

  • ATLAS reports ( raw view) -
    • Activities: Reprocessing of data16/data15 is done for the Physics_Main stream, other small streams submitted, well on schedule. MC16 production continue in full speed.
    • Problems:
      • VOMS servers host certificate renewal:
        • This was done on Friday, 5 days before they were due to expire
        • The certificate for lcg-voms2 was not renewed correctly
        • This caused Panda to reject pilot heartbeats which used proxies with extensions from lcg-voms2
        • The certificate was fixed, luckily before the heartbeat timeout caused jobs to be killed
Gavin commented that the monitoring will be checked as certificates should be refreshed automatically.

  • CMS reports ( raw view) -
    • CMS recorded first beam splashes in 2017
    • Ramping of magnet is planned for this week
    • Moderate overall CPU usage
    • CMS HLT can read from CERN EOS without GSI authentication
      • Should help to lower high (authentication) load
    • Saw some authentication issues - mainly in data transfers end of last week
      • Most caused by proxies with lifetime longer than old CMS VOMS host certificate (expires on May 3rd)
      • Host certificate got renewed on Friday
      • Freshly created proxies cured authentication problems
    • Parts of CMSweb backend developed an issue on May 1st
      • Some CMS services hosted there started to misbehave
      • Taken care of by cmsweb operators on May 2nd

  • ALICE -
    • apologies: the ALICE operations experts cannot attend today
    • very high activity on average
    • CERN: all CREAM CEs were unusable for many hours on Sat (GGUS:127997)
      • thanks for fixing that the same day!

Sites / Services round table:

  • ASGC: ntr
  • BNL: AFS phase-out from Tier1 WNs, ongoing
  • CNAF:
    • Unscheduled downtime today and tomorrow for ATLAS and LHCb storage. Reported on GOCDB. There is another T1 down. We have a hardware failure on various LUN on the same controller.
  • EGI: nc
  • FNAL: ntr
  • IN2P3: nc
  • JINR: HC at 0% at 28-04-2017 for unknown reasons. Investigating.
  • KISTI: nc
  • KIT: nc
  • NDGF: nc
  • NL-T1:
    • Since yesterday, SURFsara dCache SRM is running dual stack IPv4 & IPv6. So far we have not observed problems. We've observed some IPv6 traffic.
      • Yesterday's maintenance was submitted in GOCDB as only a warning, while it should have been an outage. Our apologies for the confusion.
  • NRC-KI: nc
  • OSG: New accounting system was put in place, it will be moved to Uni of Nebraska as next step
  • PIC: ntr
  • RAL: Lots of problems with Castor over the weekend. We are speeding up testing of Castor 2.1.16
  • TRIUMF: ntr

  • CERN computing services:
    • VOMS hostcert renewed with 5 days to go - but renewed cert had problems.
    • CREAM CEs GridFTP problem due to a bug introduced in the site gridFTP configuration.
  • CERN storage services: NTR
  • CERN databases:
    • Network DB migration tomorrow - no creation or deletion of VMs will be possible at the time (OTG:0035934)
  • GGUS: NTR
  • Monitoring:
    • Final reports for the March 2017 availability sent around
  • MW Officer:
    • VOMS host certificates renewed at CERN on Friday lunch time. The previous certs will expire tomorrow ~12:00 therefore some long lasting VOMS proxies created before Friday will start to be refused by Grid services. We suggest VOs to recreate any long lasting proxy created before Friday, in particular some of the CMS proxies delegated to FTS are affected. ( new VOMS proxy and delegation to FTS is needed). This issue comes from a VOMS bug not yet fixed ( GGUS:120463)
  • Networks:
  • Security: NTR

AOB:

Edit | Attach | Watch | Print version | History: r17 < r16 < r15 < r14 < r13 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r17 - 2017-05-02 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback